OpenStudy (anonymous):

1. The ordered pair (3, -2) is a solution to the system of inequalities y _>_ - x - 2 True or False My answer: True

OpenStudy (radar):

Was it right?

OpenStudy (radar):

\[y \ge-x-2\]Is that what is suppose to be represented above in your problem?

OpenStudy (anonymous):

perf

OpenStudy (radar):

0 -2>=-3-2 -2>= -5 That would be true wouldn't it?